OpenStudy (anonymous):

Write the geometric sequence that has 4 means between 1/2 and -1/64 ?

OpenStudy (anonymous):

Find the Common Ratio First

OpenStudy (anonymous):

how

OpenStudy (anonymous):

6 th term is -1/64

OpenStudy (anonymous):

and 1st term is 1/2

OpenStudy (anonymous):

a*r^5 = -1/32 a = 1/2.....find r?

OpenStudy (anonymous):

1/2^5 ??